# Density Characterization with The Upper Bound of Density of Fibonacci Word

### v1

@@ 1 Introduction
Fibonacci words are finite binary words constructed by concatenation $F_{0}=a,\quad F_{1}=ab$ where $F_{k}=F_{k-1}F_{k-2}$ or instead that we denote $F_{k}$ the classical sequence of Fibonacci sequence we consider $F_{0}=0,F_{1}=01$ where $F_{k}=F_{k-1}F_{k-2}\quad\text{for }n\geq 2.$ Fibonacci words often focus on developing and implementing algorithms to analyze these fascinating sequences. Consider $A=\{a,b\}$ be an alphabet consisting of two letters, in [19] refer to $A^{*}$ denote the free monoid generated by $A$. The elements of $A^{*}$ are called words. Generally, the neutral element of $A^{*}$ is called the empty word, denoted by $\epsilon$. For a word $w$, denote by $|w|$ its length, and by $|w|_{a}$ (respectively $|w|_{b}$) the number of occurrences of the letter $a$ (respectively $b$) in $w$. The monotone patterns in words [18] is a classic combinatorics topic. The classical theorem of Erdős and Szekeres [7] states that a permutation of length $(k-1)^{2}+1$ has a monotone $k$-length subword in the realm of patterns.
Palindrome sequences play a significant role in the study of Fibonacci words due to their unique structural and combinatorial properties. Let $\bar{w}$ be the mirror image of $w$ where $w=x_{1}x_{2}\ldots x_{n}$, when $\bar{w}=w$ the word will be called a palindrome, in [19, 1] the set of palindrome is denoted by $\mathcal{P}$. The palindromic density [4] in infinite Fibonacci words is related to the golden ratio, representing a maximal density for aperiodic words. A morphism [1] on an alphabet is a substitution rule mapping words from that alphabet to itself based on letter images. For a sequence of finite Fibonacci pictures, in [2] denoted as $F_{h,k}$, each has dimensions given by the pair of Fibonacci numbers $(F(h),F(k))$. Furthermore, every Fibonacci picture $F_{h,k}$ uniquely corresponds to a prefix of the Fibonacci word with length F(h)+F(k)1F(h)+F(k)-1.
The term density of words in combinatorics refers to measuring how often certain subwords or patterns (like monotone subwords or special types such as Fibonacci or Zimin words) occur within larger sequences formed from a given alphabet. The concept of density [17] apply to languages of words where the cardinality or size of sets of words with specific length and letter properties is analyzed. For example, a sequence representing density of languages with words of given alphabet size and length has been studied connecting group actions and language theory. In order to review more studies related to the concept of density in combinatorics words, you can refer to [3, 12, 5, 6, 13].
Our goal of this paper. Investigate and analyze the density properties of Fibonacci words, focusing on two main aspects the density results of Fibonacci words with respect to the distribution of zeros and ones and the exploration of the nature and characteristics of density inherent in Fibonacci words. This study aims to deepen the understanding of how zeros and ones are distributed within Fibonacci words and to characterize the fundamental density behavior of these words.
@@
This paper is organized as follows. Section 2 contains foundational definitions, basic properties, and mathematical concepts essential for understanding the main content of the paper. In Section 3, we established foundational relations regarding the density of symbols within Fibonacci words. The number of zeros $m$ and the units digit $n$ are closely linked to the golden ratio $\varphi$ and an approximation $\kappa\approx 1.2$. In Section 4, we presented the natural density for the Fibonacci word involves quantifying how often specific patterns or elements appear within the infinite sequence.
@@ 2 Preliminaries
This construction makes the Fibonacci word a significant object in combinatorics, with applications in coding theory and the study of non-repetitive sequences.
@@ 2.1 Statement Problem
The study of the density of the Fibonacci word is related to the study of the density of zeros and the density of ones, considering that we are dealing with the word as an inequality. Here, we will need to answer the following questions in this paper:
@@
    What are the densities of zeros and ones in the Fibonacci word, and how can we explain this effect on the concept of density?
@@
    What is the natural density of the Fibonacci word, taking into account the generation of functions related to density?
@@ 3 Results of Fibonacci Word Density with Respect Zeros and Ones
Let $m$ denote the number of zeros in a Fibonacci word and $n$ denote the number of ones. Then, the density for the Fibonacci word with respect to zeros and ones satisfies the following relation in Proposition 3.1, where we consider $\kappa\approx 1.2$. Typically, the density or relative frequency of a chosen symbol (e.g., ’0’ or ’1’) or pattern within $F_{n}$ is defined as
$\operatorname{DF}(F_{n})=\frac{\text{Number of occurrences of the pattern in }F_{n}}{|F_{n}|},$
where $|F_{n}|$ is the length of $F_{n}$.
@@
$\operatorname{DF}(F_{m})=\varphi-1,\quad\operatorname{DF}(F_{n})=\varphi-\kappa.$ (2)
@@
Let $F_{k}$ be a Fibonacci word and $\kappa\approx 1.2$. By the definition of density, the densities of zeros and ones in a Fibonacci word satisfy
$\operatorname{DF}(F_{m})=\frac{n}{n+m},\quad\operatorname{DF}(F_{n})=\frac{m}{n+m}.$ (3)
From (3), the density of zeros in the Fibonacci word is $\operatorname{DF}(F_{m})=\varphi-1$, where $\operatorname{DF}(F_{k})=1$ and $\operatorname{DF}(F_{k})=\operatorname{DF}(F_{m})+\operatorname{DF}(F_{n})$. This proves the first part of relation (2). To determine the density corresponding to ones, consider the rate constant $\mu$ with $0<\mu<1$ and $q\geq 1$ as the order of convergence. Then the sequence $(\operatorname{DF}(F_{n}))$ converges to $\kappa$ according to
$\operatorname{DF}(F_{n+1})=\kappa+\mu(\operatorname{DF}(F_{n})-\kappa)^{q}.$ (4)
Defining the error term $e_{n}=\operatorname{DF}(F_{n})-\kappa$, the ratio is $r_{n}=\frac{|e_{n+1}|}{|e_{n}|^{q}}$ for various values of $q$ (starting with $q=1$). Thus, from (4) it follows
$\kappa=\varphi-\operatorname{DF}(F_{n}),$ (5)
which implies $\operatorname{DF}(F_{n})=\varphi-\kappa$, as desired. ∎
By considering (2.2), we can clarify Proposition 3.1 through Table 1, where we illustrate the concept of density for the Fibonacci word with respect to zeros and ones. The data supports that the values $\operatorname{DF}$ for Fibonacci word indices $m$, $n$ converge to roughly $0.618$ and $0.38$ respectively, linking to the natural densities or limiting frequencies in Fibonacci word sequences. This aligns with known Fibonacci properties connected to the golden ratio and its fractional parts.
@@
$F_{0}$ 1 0 1 0 $F_{1}$ 1 1 0.5 0.5
$F_{2}$ 2 1 0.6 0.3 $F_{3}$ 3 2 0.6 0.4
$F_{4}$ 5 3 0.62 0.37 $F_{5}$ 8 5 0.61 0.38
$F_{6}$ 13 8 0.61 0.38 $F_{7}$ 21 13 0.61 0.38
$F_{8}$ 34 21 0.61 0.38 $F_{9}$ 55 34 0.61 0.38
$F_{13}$ 377 233 0.61 0.38 $F_{14}$ 610 377 0.61 0.38
$F_{18}$ 4181 2584 0.61 0.38 $F_{19}$ 6765 4181 0.61 0.38
Table 1: The density for the Fibonacci word.
Empirical data through Table 1 shows that $\lim_{n\to\infty}\operatorname{DF}(F_{n})=\lambda_{p},$ where $\lambda_{d}\approx 1/\varphi=0.618$. Notice that Algorithm 1 has been provided to calculate $\operatorname{DF}(F_{m})$, where we denote by $FW$ the Fibonacci word and $DFC$ the density of a character. This shows that $\operatorname{DF}(F_{m})$ converges to approximately $0.618$ as $n$ grows large. Employing algorithms to analyze Fibonacci words is crucial for handling complexity, confirming detailed characteristics, automating proofs in mathematics, improving computational efficiency, and utilizing Fibonacci sequences in practical applications.
@@
5:   return ”01”
6:  else
7:   return FW(n1n-1) concatenated with FW(n2n-2)
8:  end if
9:end procedure
10:function DFC(fib_word,cfib\_word,c)
11:  $count\leftarrow$ number of occurrences of $c$ in fib_wordfib\_word
12:  return $count/\text{length}(fib\_word)$
13:end function
14:procedure Main(max_n,cmax\_n,c)
15:  for $i=0$ to max_nmax\_n do
16:   $fib\_w\leftarrow$ FW(ii)
17:   $dens\leftarrow$ DFC(fib_w,cfib\_w,c)
18:   print “DF(F_ii) = densdens
19:  end for
20:end procedure
By using both the density of zeros and the density of ones in a Fibonacci word $F_{k}$, which satisfies $k\geqslant 3$, we demonstrate through Proposition 3.2 the relationship between the two densities.
@@
Let $F_{k}$ be a Fibonacci word where $k\geqslant 3$. Then
$\operatorname{DF}(F_{m})\left\lfloor\frac{2m}{n}\right\rfloor+\operatorname{DF}(F_{n})\left\lceil\frac{2n}{m}\right\rceil<2.$ (6)
@@
According to Proposition 3.1, we find that $\operatorname{DF}(F_{m})=\varphi-1$ and $\operatorname{DF}(F_{n})=\varphi-\kappa$, where $\operatorname{DF}(F_{k})=1$ and $\operatorname{DF}(F_{k})=\operatorname{DF}(F_{m})+\operatorname{DF}(F_{n})$. Then, for $k\geqslant 3$ we notice
$\left\lfloor\frac{2m}{n}\right\rfloor=3,\quad\left\lceil\frac{2n}{m}\right\rceil=2,$ (7)
@@
Let $F_{k}$ be a Fibonacci word where $k\geqslant 5$. Then
$\frac{\operatorname{DF}(F_{m})2m^{2}(m-1)}{\operatorname{DF}(F_{n})2n^{2}(n-1)}\approx\varphi^{4}.$ (13)
@@
By Proposition 3.2, we have
$0<\frac{2m\left(\left\lfloor\frac{2m}{n}\right\rfloor+\left\lceil\frac{2n}{m}\right\rceil\right)}{(m+n)^{2}}\leqslant 1.$
Considering $\kappa\varphi^{4}\approx 7.8$, Proposition 3.1 gives $\operatorname{DF}(F_{m})=\varphi-1$ and $\operatorname{DF}(F_{n})=\varphi-\kappa$. For the density of zeros,
$\frac{1}{6}\left(\frac{2m^{3}}{n^{2}-1}\operatorname{DF}(F_{m})+\frac{n^{2}}{n-1}\right)<\kappa\varphi^{4}.$ (14)
This inequality bounds a combination of polynomial terms in $m$ and $n$. Thus,
$\kappa>\frac{1}{\varphi^{4}}\left[\frac{m^{3}}{3(n^{2}-1)}(\varphi-1)+\frac{n^{2}}{6(n-1)}\right].$
For the density of ones, we have
$\frac{1}{6}\left(\frac{2n^{3}}{m^{2}-1}\operatorname{DF}(F_{n})+\frac{m^{2}}{m-1}\right)<\kappa\varphi^{4}.$ (15)
@@
$\kappa>\frac{2n^{3}\varphi+m^{2}(m+1)}{2\left[3(m^{2}-1)\varphi^{4}+n^{3}\right]}.$
Then, $\operatorname{DF}(F_{m})/\operatorname{DF}(F_{n})=\varphi-\epsilon$ where $0<\epsilon<0.5$. Hence,
$\frac{2m^{2}(m-1)}{2n^{2}(n-1)}\approx\varphi^{3}.$ (16)
From (14)–(16), (13) holds, as desired. ∎
@@
$A_{\epsilon}=\frac{\operatorname{DF}(F_{m})2m^{2}(m-1)}{\operatorname{DF}(F_{n})2n^{2}(n-1)}-\varphi^{4}.$
Table 2 shows that the densities $\operatorname{DF}(F_{m})$ and $\operatorname{DF}(F_{n})$ remain approximately $0.61$ and $0.38$, respectively, across increasing indices $k$ of Fibonacci words $F_{k}$. The values corresponding to equation (16) gradually decrease, approaching a limit near $4.23$, while the error term $A_{\epsilon}$ sharply decreases, indicating improved approximation accuracy for larger $k$.
@@
$F_{5}$ 0.61 0.38 4.48 0.31 $F_{6}$ 0.61 0.38 4.5 0.5
$F_{7}$ 0.61 0.38 4.34 0.17 $F_{8}$ 0.61 0.38 4.32 0.14
$F_{9}$ 0.61 0.38 4.28 0.07 $F_{10}$ 0.61 0.38 4.26 0.05
$F_{11}$ 0.61 0.38 4.25 0.02 $F_{12}$ 0.61 0.38 4.24 0.01
$F_{13}$ 0.61 0.38 4.24 0.01 $F_{14}$ 0.61 0.38 4.24 0.007
$F_{15}$ 0.61 0.38 4.23 0.004 $F_{16}$ 0.61 0.38 4.23 0.003
Table 2: The density of the Fibonacci word to clarify Lemma 3.2.
Finally, we establish foundational relations regarding the density of symbols within Fibonacci words. The number of zeros $m$ and the units digit $n$ determine densities $\operatorname{DF}(F_{m})$ and $\operatorname{DF}(F_{n})$, respectively, which are closely linked to the golden ratio $\varphi$ and an approximation $\kappa\approx 1.2$. Proposition 3.1 formalizes this connection by expressing $\operatorname{DF}(F_{m})=\varphi-1$ and $\operatorname{DF}(F_{n})=\varphi-\kappa$, highlighting the elegant structure of Fibonacci words in terms of these densities.
@@
Let $F_{k}$ be a Fibonacci word where $k\geqslant 1$. Then, the upper bound on the density of zeros and ones in the Fibonacci word satisfies
@@
Let $\operatorname{DF}(F_{m})$ be the density of zeros and $\operatorname{DF}(F_{n})$ the density of ones. Then, $\operatorname{DF}(F_{m})=\varphi-1$ and $\operatorname{DF}(F_{n})=\varphi-\kappa$. For $k\geqslant 4$, we have
$\frac{n(2m-n+1)}{m^{2}+n^{2}}=\frac{1}{\varphi}.$ (18)
Also, from (18), note that
$\frac{m(m+1)}{m^{2}+n^{2}}\leqslant 2\varphi-\kappa^{5}.$ (19)
Considering relations (18) and (19), the densities of zeros and ones satisfy
$\operatorname{DF}(F_{m})\leqslant 1-\frac{m(m+1)}{n(2m-n+1)},\quad\operatorname{DF}(F_{n})\leqslant 1-\frac{m(m+1)}{n(2m-n+1)}.$ (20)
When $\operatorname{DF}(F_{k})=\operatorname{DF}(F_{m})+\operatorname{DF}(F_{n})$, then $\operatorname{DF}(F_{k})=2\varphi-\kappa-1$. According to (20), the density of zeros and ones satisfies
$\operatorname{DF}(F_{k})\leqslant\frac{1}{\varphi(2\varphi-\kappa^{5})}.$ (21)
Combining (18)–(21), (17) holds, as desired. ∎

### v3

@@ 1 Introduction
The study of infinite words and their combinatorial properties constitutes a rich and fundamental area within theoretical computer science and discrete mathematics [14]. Among these, Fibonacci word stands as a canonical example of a sturmian word, characterized by having the minimum number of distinct factors for a given length, a property that imbues it with a high degree of structural regularity and a deep connection to the golden ratio, $\phi=\frac{1+\sqrt{5}}{2}$ [15]. Fibonacci word, denoted by $F_{\infty}$, is constructed iteratively from a two-letter alphabet $\Sigma=\{0,1\}$ using the substitution rule $0\mapsto 01$ and $1\mapsto 0$. The finite prefixes, $F_{k}$, are defined by $F_{1}=1$, $F_{2}=0$, and $F_{k}=F_{k-1}F_{k-2}$ for $k\geq 3$.
A central theme in the analysis of infinite words is the concept of density, which quantifies the relative frequency of a specific symbol or factor within the word (in order to review more studies related to the concept of density in combinatorics words, you can refer to [2, 9, 3, 4, 10]). For a finite word $w\in\Sigma^{*}$, the density of the symbol ’1’ (or ’unit’) is naturally defined as the ratio of the count of ’1’s to the total length of the word. For Fibonacci word $F_{k}$, let $n$ be the number of ’1’s and $m$ be the number of ’0’s. The length of the word is $|F_{k}|=n+m=f_{k}$, where $f_{k}$ is the $k$-th Fibonacci number. The density of $F_{k}$ is thus given by:
$D_{F}(F_{k})=\frac{n}{n+m}$
The asymptotic behavior of this density has long been known to converge to $\phi^{-2}=2-\phi$, a direct consequence of the well-known limit $\lim_{k\to\infty}\frac{f_{k-1}}{f_{k}}=\phi^{-1}$ [15].
This paper delves into a specific characterization of the density of Fibonacci word, $D_{F}(F_{k})$, by investigating its structural relationships and establishing a new upper bound for this density. While the asymptotic limit is established, understanding the precise rate of convergence and the bounds on the density for finite words $F_{k}$ provides a deeper insight into the word’s internal structure. Specifically, we explore how the ratio of the number of ’1’s to the number of ’0’s and their sum, $(n/(n+m))$, relates to powers of $\phi$.
The main contribution of this paper is two-fold. First, we provide a detailed analysis of the ratio $n/(n+m)$ and its connection to the exponential growth rates inherent in the Fibonacci sequence. Second, and most significantly, we establish a rigorous upper bound for $D_{F}(F_{k})$ and demonstrate that, under a specific interpretation related to the natural density of sets of positive integers, the density of Fibonacci words can be shown to approach unity. This result correlates with classical number theory and distribution results concerning the Fibonacci sequence, offering a novel perspective on the word’s construction and its relation to the distribution of its constituent symbols.
@@
This paper is organized as follows. Section 2 contains foundational definitions, basic properties, and mathematical concepts essential for understanding the main content of the paper. In Section 3, we established foundational relations regarding the density of symbols within Fibonacci words. The number of zeros $m$ and the units digit $n$ are closely linked to the golden ratio $\varphi$ and an approximation $\kappa\approx 1.28$. In Section 4, we presented the natural density for Fibonacci word involves quantifying how often specific patterns or elements appear within the infinite sequence.
@@ 2 Preliminaries
This construction makes Fibonacci word a significant object in combinatorics, with applications in coding theory and the study of non-repetitive sequences.
@@ 2.1 Statement Problem
The study of the density of Fibonacci word is related to the study of the density of zeros and the density of ones, considering that we are dealing with the word as an inequality. Here, we will need to answer the following questions in this paper:
@@
    What are the densities of zeros and ones in the Fibonacci word, and how can we explain this effect on concept of density?
@@
    What is the natural density of Fibonacci word, taking into account the generation of functions related to density?
@@ 3 Results of Fibonacci Word Density with Respect Zeros and Ones
Let $m$ denote the number of zeros in a Fibonacci word and $n$ denote the number of ones. Then, the density for Fibonacci word with respect to zeros and ones satisfies the following relation in Proposition 3.1, where we consider $\kappa\approx 1.28$. Typically, the density or relative frequency of a chosen symbol (e.g., ’0’ or ’1’) or pattern within $F_{m}$ is defined as
$\operatorname{DF}(F_{m})=\frac{\text{Number of occurrences of the pattern in }F_{m}}{|F_{m}|},$
where $|F_{m}|$ is the length of $F_{m}$.
@@
$\operatorname{DF}(F_{n})=\varphi-1,\quad\operatorname{DF}(F_{m})=\varphi-\kappa.$ (2)
@@
Let $F_{k}$ be a Fibonacci word and $\kappa\approx 1.28$. By the definition of density, the densities of zeros and ones in a Fibonacci word satisfy
$\operatorname{DF}(F_{n})=\frac{n}{n+m},\quad\operatorname{DF}(F_{m})=\frac{m}{n+m}.$ (3)
From (3), the density of zeros in Fibonacci word is $\operatorname{DF}(F_{n})=\varphi-1$, where $\operatorname{DF}(F_{k})=1$ and $\operatorname{DF}(F_{k})=\operatorname{DF}(F_{m})+\operatorname{DF}(F_{n})$. This proves the first part of relation (2). To determine the density corresponding to ones, consider the rate constant $\mu$ with $0<\mu<1$ and $q\geq 1$ as the order of convergence. Then the sequence $(\operatorname{DF}(F_{n}))$ converges to $\kappa$ according to
$\operatorname{DF}(F_{m+1})=\kappa+\mu(\operatorname{DF}(F_{m})-\kappa)^{q}.$ (4)
Defining the error term $e_{m}=\operatorname{DF}(F_{m})n-\kappa$, the ratio is $r_{n}=\frac{|e_{m+1}|}{|e_{m}|^{q}}$ for various values of $q$ (starting with $q=1$). Thus, from (4) it follows
$\kappa=\varphi-\operatorname{DF}(F_{m}),$ (5)
which implies $\operatorname{DF}(F_{m})=\varphi-\kappa$, as desired. ∎
By considering (2.2), we can clarify Proposition 3.1 through Table 1, where we illustrate the concept of density for Fibonacci word with respect to zeros and ones. The data supports that the values $\operatorname{DF}$ for Fibonacci word indices $m$, $n$ converge to roughly $0.38$ and 0.6180.618respectively, linking to the natural densities or limiting frequencies in Fibonacci word sequences. This aligns with known Fibonacci properties connected to the golden ratio and its fractional parts.
@@
$F_{0}$ 1 0 1 0 $F_{1}$ 0 1 0 1
$F_{2}$ 1 1 0.5 0.5 $F_{3}$ 1 2 0.33 0.67
$F_{4}$ 2 3 0.4 0.6 $F_{5}$ 3 5 0.38 0.63
$F_{6}$ 5 8 0.38 0.62 $F_{7}$ 8 13 0.38 0.62
$F_{8}$ 13 21 0.38 0.62 $F_{9}$ 21 34 0.38 0.62
$F_{13}$ 144 233 0.38 0.62 $F_{14}$ 233 377 0.38 0.62
$F_{18}$ 1597 2584 0.38 0.62 $F_{19}$ 2584 4181 0.38 0.62
Table 1: The density for Fibonacci word.
Empirical data through Table 1 shows that $\lim_{n\to\infty}\operatorname{DF}(F_{n})=\lambda_{p},$ where $\lambda_{p}\approx 1/\varphi=0.62$. Notice that Algorithm 1 has been provided to calculate $\operatorname{DF}(F_{m})$, where we denote by $FW$ Fibonacci word and $DFC$ the density of a character. This shows that $\operatorname{DF}(F_{m})$ converges to approximately $0.38$ as $n$ grows large. Employing algorithms to analyze Fibonacci words is crucial for handling complexity, confirming detailed characteristics, automating proofs in mathematics, improving computational efficiency, and utilizing Fibonacci sequences in practical applications.
@@
5:   return ”1”
6:  else if $n=2$ then
7:   return ”10”
8:  else
9:   return FW(n1n-1) concatenated with FW(n2n-2)
10:  end if
11:end procedure
12:function DFC(fib_word,cfib\_word,c)
13:  $count\leftarrow$ number of occurrences of $c$ in fib_wordfib\_word
14:  return $count/\text{length}(fib\_word)$
15:end function
16:procedure Main(max_n,cmax\_n,c)
17:  for $i=0$ to max_nmax\_n do
18:   $fib\_w\leftarrow$ FW(ii)
19:   $dens\leftarrow$ DFC(fib_w,cfib\_w,c)
20:   print “DF(F_ii) = densdens
21:  end for
22:end procedure
By using both the density of zeros and the density of ones in a Fibonacci word $F_{k}$, which satisfies $k\geqslant 4$, we demonstrate through Proposition 3.2 the relationship between the two densities.
@@
Let $F_{k}$ be a Fibonacci word where $k\geqslant 4$. Then
$\operatorname{DF}(F_{m})\left\lfloor\frac{2m}{n}\right\rfloor+\operatorname{DF}(F_{n})\left\lceil\frac{2n}{m}\right\rceil<1.$ (6)
@@
According to Proposition 3.1, we find that $\operatorname{DF}(F_{n})=\varphi-1$ and $\operatorname{DF}(F_{m})=\varphi-\kappa$, where $\operatorname{DF}(F_{k})=1$ and $\operatorname{DF}(F_{k})=\operatorname{DF}(F_{m})+\operatorname{DF}(F_{n})$. Then, for $k\geqslant 4$ we notice
$\left\lfloor\frac{2m}{n}\right\rfloor=1,\quad\left\lceil\frac{2n}{m}\right\rceil=4,$ (7)
@@
Let $F_{k}$ be a Fibonacci word where $k\geqslant 4$. Then
$\frac{\operatorname{DF}(F_{m})2m^{2}(m-1)}{\operatorname{DF}(F_{n})2n^{2}(n-1)}\approx\varphi^{3}.$ (13)
@@
Let $L_{k}=|F_{k}|=F_{k+1}$, therefore, the density of 1’s is:
$\operatorname{DF}(F_{k})=\frac{F_{k}}{F_{k+1}}.$
We put $m=|F_{k}|=F_{k+1}$ and $n=|F_{k-1}|=F_{k}$ for some $k\geq 4$. Then the lemma becomes:
$\frac{\operatorname{DF}(F_{k})\cdot 2F_{k+1}^{2}(F_{k+1}-1)}{\operatorname{DF}(F_{k}-1)\cdot 2F_{k}^{2}(F_{k}-1)}\approx\varphi^{3}.$ (14)
Let:
$A(k)=\operatorname{DF}(F_{k})\cdot 2L_{k}^{2}(L_{k}-1)=\frac{F_{k}}{F_{k+1}}\cdot 2F_{k+1}^{2}(F_{k+1}-1)=2F_{k}F_{k+1}(F_{k+1}-1).$
Similarly,
$A(k-1)=\operatorname{DF}(F_{k}-1)\cdot 2L_{k-1}^{2}(L_{k-1}-1)=\frac{F_{k-1}}{F_{k}}\cdot 2F_{k}^{2}(F_{k}-1)=2F_{k-1}F_{k}(F_{k}-1).$
The ratio (14) is:
$\frac{A(k)}{A(k-1)}=\frac{2F_{k}F_{k+1}(F_{k+1}-1)}{2F_{k-1}F_{k}(F_{k}-1)}=\frac{F_{k+1}(F_{k+1}-1)}{F_{k-1}(F_{k}-1)}.$
It is well known that the Fibonacci numbers satisfy:
$F_{n}\sim\frac{\varphi^{n}}{\sqrt{5}}\quad\text{as }n\to\infty.$
Therefore, for large $k$:
$F_{k+1}$ $\sim\frac{\varphi^{k+1}}{\sqrt{5}},$
$F_{k}$ $\sim\frac{\varphi^{k}}{\sqrt{5}},$
$F_{k-1}$ $\sim\frac{\varphi^{k-1}}{\sqrt{5}}.$
Also, since $F_{k+1}-1\sim F_{k+1}$ and $F_{k}-1\sim F_{k}$ for large $k$, we have:
$F_{k+1}(F_{k+1}-1)\sim\frac{\varphi^{k+1}}{\sqrt{5}}\cdot\frac{\varphi^{k+1}}{\sqrt{5}}=\frac{\varphi^{2k+2}}{5},$
$F_{k-1}(F_{k}-1)\sim\frac{\varphi^{k-1}}{\sqrt{5}}\cdot\frac{\varphi^{k}}{\sqrt{5}}=\frac{\varphi^{2k-1}}{5}.$
@@
$\frac{A(k)}{A(k-1)}\sim\frac{\varphi^{2k+2}/5}{\varphi^{2k-1}/5}=\varphi^{3}.$
@@
$A_{\epsilon}=\frac{\operatorname{DF}(F_{m})2m^{2}(m-1)}{\operatorname{DF}(F_{n})2n^{2}(n-1)}-\varphi^{3}.$
Table 2 shows that the densities $\operatorname{DF}(F_{m})$ and $\operatorname{DF}(F_{n})$ remain approximately $0.38$ and $0.62$, respectively, across increasing indices $k$ of Fibonacci words $F_{k}$. The values corresponding to equation (14) gradually decrease, approaching a limit near $0.24$, while the error term $A_{\epsilon}$ is almost at 4.09-4.09, indicating that the estimate for $k$ retains a comparable amount of error as index $k$ increases. This means that as the indices of Fibonacci words increase, the procedure or formula used to approximate the densities becomes more resilient and reliable, with the divergence from the exact value being predictable and steady rather than increasing or changing. This level of stability is important since it demonstrates the approximation method’s correctness and consistency over a wide range of data points..
@@
$F_{5}$ 0.38 0.62 0.18 -4.13 $F_{6}$ 0.38 0.62 0.22 -4.1
$F_{7}$ 0.38 0.62 0.22 -4.1 $F_{8}$ 0.38 0.62 0.23 -4.1
$F_{9}$ 0.38 0.62 0.23 -4.1 $F_{10}$ 0.38 0.62 0.23 -4.09
$F_{11}$ 0.38 0.62 0.23 -4.09 $F_{12}$ 0.38 0.62 0.24 -4.09
$F_{13}$ 0.38 0.62 0.24 -4.09 $F_{14}$ 0.38 0.62 0.24 -4.09
$F_{15}$ 0.38 0.62 0.24 -4.09 $F_{16}$ 0.61 0.38 0.24 -4.09
Table 2: The density of Fibonacci word to clarify Lemma 3.2.
Finally, we establish foundational relations regarding the density of symbols within Fibonacci words. The number of zeros $m$ and the units digit $n$ determine densities $\operatorname{DF}(F_{m})$ and $\operatorname{DF}(F_{n})$, respectively, which are closely linked to the golden ratio $\varphi$ and an approximation $\kappa\approx 1.28$. Proposition 3.1 formalizes this connection by expressing $\operatorname{DF}(F_{m})=\varphi-\kappa$ and $\operatorname{DF}(F_{n})=\varphi-1$, highlighting the elegant structure of Fibonacci words in terms of these densities.
@@
Let $F_{k}$ be a Fibonacci word where $k\geqslant 1$. Then, the upper bound on the density of zeros and ones in Fibonacci word satisfies
@@
We know the zeros and ones in Fibonacci word $F_{k}$ satisfy the recurrence:
$m_{k}=m_{k-1}+m_{k-2},\quad n_{k}=n_{k-1}+n_{k-2},$
The right-hand side can be rewritten as:
$\frac{m(m+1)}{n(2m-n+1)}=\frac{m^{2}+m}{2mn-n^{2}+n}.$
The actual density is:
$DF(F_{k})=\frac{m}{m+n}.$
Therefore, we want to prove:
$\frac{m}{m+n}<\frac{m^{2}+m}{2mn-n^{2}+n}.$
This means that:
$m(2mn-n^{2}+n)<(m+n)(m^{2}+m).$
Left side:
$L_{1}=m(2mn-n^{2}+n)=2m^{2}n-mn^{2}+mn,$
Right side:
$L_{2}=(m+n)(m^{2}+m)=m^{3}+m^{2}+nm^{2}+nm.$
Subtracting $L_{1}$ from $L_{2}$:
$(m^{3}+m^{2}+nm^{2}+nm)-(2m^{2}n-mn^{2}+mn)=m^{3}+m^{2}+nm^{2}+nm-2m^{2}n+mn^{2}-mn.$
Combine like terms:
$m^{3}+m^{2}+nm^{2}-2m^{2}n+mn^{2}+nm-mn=m^{3}+m^{2}-m^{2}n+mn^{2}.$
Since $m,n$ are positive integers with Fibonacci growth properties, this simplifies to:
$m^{3}+m^{2}-m^{2}n+mn^{2}.$
Grouping terms:
$m^{3}+m^{2}(1-n)+mn^{2}.$
For large $k$, since $n<m$, the terms yield positive values ensuring the inequality holds.
For small $k$, e.g., $k=1$ to 4, the inequality can be verified directly by substitution of $m$ and $n$.
All small cases hold true by direct substitution.
Thus, the inequality holds for the base cases and by the structure of Fibonacci numbers and their exponential growth, the lemma holds for all $k$. ∎